OpenStudy (anonymous):

How many solutions does the equation 4sin2(2x) -1 = 0 have in the interval [0, 2Ï€]?

OpenStudy (anonymous):

it is the same ques.

OpenStudy (anonymous):

sin^2(2x)=1/4

OpenStudy (anonymous):

so there's only one solution?

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

what sol. you got?

OpenStudy (anonymous):

\[2x=n \pi \pm \alpha\] \[x=\frac{n \pi \pm \alpha}{2}\]

OpenStudy (anonymous):

now put values of n till your ans. is less than 2 pi.